# Fast In-Spectrum Graph Watermarks

### v1

@@ IV-F1 Intensity and Impact of the Attack
In order to assess a plausible intensity of an attack that a graph can sustain without losing its utility, we show in Figure 4 a simple function performed on the three first graphs of Table II: the ranking of the top-$1000$ nodes according to the degree centrality [6]. The figure plots the changes in this ranking of the most central nodes, according to the classic Spearman metric [10]. It measures the correlation between two rankings of the top nodes (on the original graph and on the attacked version). The correlation drops dramatically as the attack increases, with even the low value of $1\%$ of edges flipped. Thus, it appears that this random attack disrupts the connectivity of the central nodes, clearly questioning the usefulness of the resulting graph. In this context, $10\%$ of flipped edges is an extreme attack that appears to remove the utility of the graph, and can thus be considered a very conservative upper bound.
@@
IV-F2 Robustness
The performance of FFG is reported on real large graphs in Table II.
The key lengths are similar to those used in Eppstein et al. [13] and standard deviations are set to target watermarks that cause little distortion. Despite their various densities, all of them were watermarked with indeed small distortions (EDs of $10^{-6}$ at the very least, most often at $10^{-8}$, and even as low as $10^{-10}$ for the graph ”inf-belgium-osm”).
Graphs $|V|$ $\frac{|E|}{|V|}$ $m$ $\sigma$ ED $\theta$ $Suitability$
inf-belgium-osm 1.4M 1.1 54 250 $3.10^{-10}$ 0.6
soc-YouTube-ASU 1.1M 2.6 200 12,000 $1.1.10^{-8}$ 2
hollywood-2009 1.1M 52.7 162 224,000 $1.6.10^{-8}$ 205
rgg-n-2-20-s0 1M 6.6 119 1,624 $3.8.10^{-8}$ 10.2
kron-g500-logn20 1M 42.6 71128 16,000 $3.5.10^{-8}$ 374
scale21-ef16-adj 1.2 M 51 44,743 12,000 $1.7.10^{-8}$ 233
roadNet-PA 1.1M 1.4 48 184 $6.10^{-6}$ 0.6
delaunay-n20 1M 3.0 63 61 $4.3.10^{-8}$ 6.2
TABLE II: Experiment settings and results for FFG, on real large graphs.
We now subject these graphs to attacks and observe the resilience of FFG. The last column $Suitability$ shows that the extractions are successful despite the attack. We also note that additional tests have shown that all watermarks can as well be extracted even under an attack that flips $100$% edges, for a value of $\theta$ set to be resistant to only $10$ edge flips. This underscores the robustness of FFG, and the reach of Goal 4.
This experimental section confirms both the applicability of FFG to real graphs, the low distortion of the scheme on these graph structures, and its resilience to significant attack strengths.
@@ IV-F3 Densities and the Resulting Threshold
The last experiment for FFG is to test its robustness to different graph densities, for the three graph models. We take the three graph generators (BA, WS, ER) and vary the density $\frac{|E|}{N}$ of graphs to $5$, $30$ and $50$. $5$ is a low density which respects the uniqueness condition, $30$ and $50$ are medium and high densities according to [26] and the graphs taken from SNAP [24]. For each generator and each density, we set $\theta$ to be resistant to $x$edge flips ($x$-axis). This resilience is observed on $3$ runs per point on Figure 5, where $\theta$ is set by a dichotomous search for the effective resilience to the attack.
Figure 5: The choice of $\theta$ in FFG to be resistant to $x$edge flips depending on the density and the type of graph model.
Figure 5 shows that the resulting threshold $\theta$ depends primarily on the type of graph. For BA or WS graphs, a setting $\theta=1$ is sufficient for FFG to be resistant to attacks with up to $10\%$ edge flips ($x$-axis), regardless of the density of the original graph. Density is the most relevant parameter when watermarking ER graphs, where $\theta$ can vary from $0.3$ to $5$ to be resistant to attacks of up to 10% edge flips.
@@ V A Comparative Benchmark
Finally, we perform a head-to-head comparative benchmark of FFG with the two state-of-the-art approaches [38] (Zhao) and [13] (Eppstein), both in terms of running time and resilience to attacks.
@@ V-A Related work
There are four steps in the graph embedding function of Zhao et al. [38]. First, they generate a random seed based on cryptographic keys. Then, they generate a random watermark graph (Erdős-Rényi) using the seed and match it with a subgraph of $G$. The subgraph selection is based on the structure of the vertices (i.e. their neighborhoods). Finally, the embedding is performed by an XOR operation between the subgraph and the random watermark graph. For the extraction, they regenerate the watermark graph and the subgraph like in the embedding process. After that, they identify in the potential $G^{*}$ which vertices are candidates to match the subgraph (again using the vertex structure). Finally, they extract the watermark using a recursive algorithm. Since the code was not provided by the authors, we implemented it in Python 3. We note that our execution results are consistent with those presented by the authors in their experimental section (Figure 2b in [38]). We set $\theta=0.95$ and $B=25$, for an acceptable execution time.
@@
The scheme of Eppstein et al.[13] consists of the same three functions Keygen, Embed and Extract as the FFG scheme. In the scheme of Eppstein et al., private keys generated by the Keygen function are sets of vertex pairs in a graph containing as many vertices as in the graph to be watermarked. There are three main steps in the embedding function. The first step is ordering and labeling the vertices of the graph, considering degrees and using bit vectors. The vertices are divided into three categories: high, medium, and low degree vertices. Then, each edge contained in the private key is randomly flipped according to the probability of its statistical existence in the graph. The function Extract uses graph isomorphism to identify the watermarked graph. Unlike FFG and Zhao [38], this method identifies a graph between multiple (simultaneously created) watermark copies, not just a single watermark graph with a given key. The selected copy is the closest copy considering the Hamming distance.
@@ V-B Execution Timings
We first benchmark the runtime of the Embed and Extract functions of the competing watermarking schemes. The keys are generated according to the recommendations in each paper; their timings are not displayed, as they are all very fast. Scalability improvements suggested by each paper have been implemented.
@@
Figure 6: Benchmark comparison between FFG, Eppstein [13], and Zhao [38]. (a) Timing ($y$-axis) of the Embed and Extract functions for the three competing schemes, depending on the BA graph size $N$ ($x$-axis). Executions taking less than the timeout (fixed at 25 minutes) are kept and plotted. (b) Robustness of the three schemes ($y$-axis) facing attacks: the normalized number of successful extractions, facing increasing edge flips (measured by edit distance on the $x$-axis as ratio of edges) on three large graphs.
In Figure 6a, timings are measured in seconds as a function of the number of vertices $N$ in an increasing size BA graph (from $1k$ to $10M$ vertices, with an attachment parameter of $3$). We set a timeout of $25$ minutes for the experiments (dotted gray horizontal line), so that all scheme executions –each running on a single core– are discarded if they exceed this threshold. For small and medium-sized graphs ($N\leq 10^{5}$), the embedding function of Zhao et al. is the most efficient method, while its extraction takes the longest. For medium and large graphs (more than $10^{5}$ vertices), the FFG scheme outperforms the state-of-the-art. Without exceeding the timeout, FFG can watermark graphs with $10$ times more vertices than the scheme of Eppstein et al. and up to $100$ times more vertices than the scheme of Zhao et al.
Note that extractions are performed on a non-attacked $G_{W}$, which is the most favorable scenario for the schemes of Eppstein et al. and Zhao et al.: their extractions are longer when applied to attacked graphs, because the approximate subgraph matching is harder (while the matrix representation in FFG makes it time-invariant to attacks). The extractions of Zhao et al. and Eppstein et al. are linear in the number of vertices, while FFG is almost constant after $10k$ vertices.
Even though Zhao et al. and Eppstein et al. also perform dimensionality reduction, they use approximate subgraph isomorphism in the extraction, which is computationally expensive, if not intractable (NP-complete) for increasingly large graphs. FFG relies on vertex labels (in parallel with images) to avoid such a tedious operation and then works significantly faster.
@@ V-C Robustness to the Edge Flip Attack
In Table II, we observed that FFG is resistant to $100$% edge flips on real graphs. (Recall that $100\%$ is relative to the number of edges in the original graph, which here means that $|E|$ edges were flipped among the $N(N-1)/2$ possible flips.) Since Zhao and Eppstein schemes can at best achieve this resilience, we now test the watermark extractions of the three schemes, despite attacks on three watermarked graphs.
Figure 6b shows the results for the three large graphs. The success rate is shown on the $y$-axis. The $x$-axis shows the number of edges flipped, measured by the ED. Each experience is run three times (the standard deviation is zero, as there is no variation in the results across all the performed executions).
We observe that FFG performs better than the scheme of Zhao et al. for the three graphs. The scheme of Zhao et al. fails before $10$% edges are flipped, while FFG handles flips equivalent to $100\%$ edges. FFG also beats the Eppstein scheme, except on the Flickr graph where Eppstein survives $1000\%$ of flips.
We conclude that FFG is at least as robust –except on one configuration–, if not better, against random edge flips than its two competitors. Overall, the high robustness of FFG (resilience $\geq 100\%$ of flipped edges, see Table II and Figure 6b) makes it practical even in scenarios with extreme attack strengths.
@@ VI Conclusion
Large and real graphs are valuable assets that can be tracked for provenance using watermarking techniques. This paper bridges the gap between the image processing and the graph domains. While previous state-of-the-art schemes relied on standard graph manipulations, we have shown that knowledge and techniques from the multimedia community can be ported to the graph domain, yielding significant complexity improvements. This is achieved by studying the transitions between a graph and its real-valued image-like representation, and vice versa. Our method FFG shows strong performance on several metrics, including low false positives and negatives.
For future work, our framework can be extended with other binarization methods, such as the optimized thresholding method proposed by Otsu [27]. This may lead to even smaller watermarks, which is crucial for minimizing distortions in the watermarked graph. Another promising direction is to adapt recent image watermarking schemes to our framework.
Appendix A Notations
Table III is a summary of the main notations used in our paper. In particular, the differences between $\omega$, $\mathcal{W}$ and $W$ are noted. It is complemented by Figure 7, which illustrates our method.
@@ Appendix A Notations
Figure 7: The FFG framework for watermarking unweighted graphs using an image domain scheme (hereby Cox et al. [4]).
Appendix B Watermarking non-graph objects
@@
This paper is dedicated to watermark unweighted graphs. Recently, two graph watermarking approaches have been introduced that use purely graph techniques [13, 38] to watermark graphs. Since they are direct competitors to FFG, both schemes are described in detail in the main part of the paper (Section V), and then benchmarked.
@@
Appendix C Theoretical Analysis of Uniqueness
This section aims to complete the experimental study of FFG. In particular, with respect to the uniqueness and success of a watermark.
We have defined watermarking schemes and how Cox’s watermarking scheme can be adapted to graphs. We now study the effect of binarization, which is related to design goal 2.
C-A Uniqueness of Watermarks
@@
The only source of randomness in the scheme is the generation of the key; thus if $G$ is watermarked with another key $\omega_{2}$ resulting in $G_{W_{2}}$, then there is uniqueness if $G_{W_{1}}\neq G_{W_{2}}$. This property is not trivial because of the binarization which spreads keys over all the Fourier coefficients of the adjacency matrices.
We then study the probability of uniqueness:
@@
$=P\left(\bigcap_{i,j}\begin{matrix}[A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=A[i,j]=0\\ \cup A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=A[i,j]=1\\ \cup A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=1-A[i,j]=0\\ \cup A_{W_{1}}[i,j]=A_{W_{2}}[i,j]=1-A[i,j]=1]\end{matrix}\right)$
@@
Here the notation $A^{\prime}_{1,2}$ means that the result must be true for both $A^{\prime}_{1}$ and $A^{\prime}_{2}$ respectively and $av(A)$ is the average of all elements of the matrix $A$.
By the linearity of the inverse Fourier transform,
for $n\in\{1,2\}$:
$A^{\prime}_{n}=A+FT^{-1}(\mathcal{W}_{n}).$ (3)
with
$FT(A)[i,j]=\sum_{k=1}^{N}\sum_{l=1}^{N}A[k,l]e^{-2\imath\pi\left(\frac{ik}{N}+\frac{jl}{N}\right)}$
and
$FT^{-1}(A)[i,j]=\frac{1}{N^{2}}\sum_{k=1}^{N}\sum_{l=1}^{N}A[k,l]e^{2\imath\pi\left(\frac{ik}{N}+\frac{jl}{N}\right)}.$
By construction, all the elements of each key $\omega$ (and its matrix version $\mathcal{W}$) are independent. The probability $P(G_{W_{1}}=G_{W_{2}})$ in Equation (1) can be computed with the cumulative distribution of $FT^{-1}(\mathcal{W}_{n})$ ($n\in\{1,2\}$) in Equation (3).
C-B Computing the Cumulative Distribution
First, note that:
$FT^{-1}(\mathcal{W}_{n})=Z\mathcal{W}_{n}Z,$
where $Z[k,l]=\frac{1}{N}e^{2\imath\pi kl/N}$. This shows that $FT^{-1}(\mathcal{W}_{n})$ is a matrix version of a Gaussian vector. With this, we can get an exact formula for the probability distribution function of this inverse Fourier transform flattened to one dimension [15]. In fact, let $\tilde{\mathcal{W}_{n}}$ be the flattened vector of $\mathcal{W}_{n}$, where:
$\forall k\in\{1,...,N^{2}\},\tilde{\mathcal{W}_{n}}[k]=\mathcal{W}_{n}[\lceil k/N\rceil,(k-1)\mod N+1],$
where $\lceil k/N\rceil$ being the upper integer part of $k/N$ and $(k-1)\mod N$ being the rest of the Euclidean division of $k-1$ by $N$. Using these notations, we easily show that:
$FT^{-1}(\tilde{\mathcal{W}_{n}})=\tilde{Z}\mathcal{W}_{n},$
where $\tilde{Z}=Z\bigotimes Z$ is the $N^{2}\times N^{2}$ Kronecker product of $Z$ with itself. Let $(z_{k,l})$ be the coefficients of $Z$, then for $(k^{\prime},l^{\prime})\in\{1,...,N^{2}\}$:
$(\tilde{Z})_{k^{\prime},l^{\prime}}=z_{\lceil k^{\prime}/N\rceil,\lceil l^{\prime}/N\rceil}z_{(k^{\prime}-1)\mod N+1,(l^{\prime}-1)\mod N+1}.$
Since $\tilde{\mathcal{W}_{n}}$ is a Gaussian vector, $FT^{-1}(\tilde{\mathcal{W}_{n}})$ is also a Gaussian vector. Its average and standard deviation are computed using the matrix $\tilde{Z}$. Therefore, we can use standard approximation techniques to approximate Equation (C-A).
This section has successfully derived a precise analytical form for evaluating the probability of a watermark’s uniqueness. However, considering the size of the vectors involved, large graphs will require a significant approximation time. In practice, direct experiments such as those in section IV-D will be preferred. Despite this computational challenge, the analytical insights shed light on the theoretical components at stake in discretizing a watermark in an unweighted graph.
@@ C-B Computing the Cumulative Distribution
The connection between graph densities and collisions is given by the theoretical analysis in Equation (C-A).
Indeed, as $av(A)=\frac{|E|}{N^{2}}$, the threshold for binarization is equal to the density divided by $N$. For the $N=1M$ graphs (density less than $10$) in Figure 2 and the given keys:
$av(A)\simeq 0$
$avg(|FT^{-1}(W)|)\simeq 0$
$avg(|A+FT^{-1}(W)[i,j]|)\simeq av(A)$
Then, it is clear that the denser a graph, the more $A[i,j]$ coefficients are equal to $1$, and decreases the probability of collision (as seen in Equation (C-A)).
C-C Successful Embeddings: a Corollary of Uniqueness
In this section, we show that studying a successful embedding is a corollary of the analytical form of the uniqueness. The embedding of a key $\omega$ in a graph $G$ is successful if the watermarked graph $G_{W}$ is different from the original graph $G$. In other words, $G_{W}$ must be related to its key $\omega$. In fact, the key is binarized after spreading over the entire adjacency matrix. It can happen that the binarization discards a key built with insufficiently large parameters, so that no watermark is inserted into the graph. The study of successful embedding is therefore the estimation of $P(G_{W}=G)$.
@@ C-C Successful Embeddings: a Corollary of Uniqueness
Since $\texttt{Keygen}(0,0)=[]$ (an empty key) and $\texttt{Embed}(G,[])=G$, studying $P(G_{W}=G)$ is a special case of Equation (1) where $G_{W_{1}}=G_{W}$ (the graph watermarked with the key $\omega$) and $G_{W_{2}}=G$ (the graph watermarked with the empty key [][][ ]).
Note that in practice, the owner of the graph knows when the embedding has succeeded or failed: if the watermarked graph is equal to the original graph after an embedding attempt (which is linear to check in $O(|E|)$ with the edit distance for example), it is sufficient for the owner to retry an embedding operation with larger key parameters until the two graphs are truly different. (This will be studied empirically in Section IV-C.)
Appendix D Scalability reduction
In Section IV-B1, we explain that a scability reduction is applied to FFG as it is the case in  [13, 38]. This section provides the details and explanation of such an operation.
When considering graphs, the most significant information is contained in the vertices of the highest degree vertices; thus, analogous to low frequencies, in FFG, the watermark is inserted into the lowest amplitude coefficients in the spectrum among the higher-degree vertices. While the size of images on which the Cox et al. scheme is typically applied is in the order of tens of thousands of pixels, we also choose to take the top $N_{0}=10^{4}$ higher-degree vertices of $G$ to apply FFG onto.
In this way, Embed and Extract are applied only on $N_{0}$ vertices and not on the $N$ vertices of the original graph. Their worst-case time complexities, computed in Section III-D, are now $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. However, to determine the subgraph on which to apply FFG, it is necessary to select the $N_{0}$ vertices of highest degrees (which is calculated in $\mathcal{O}\left(N\log N_{0}\right)$). This cost is added naturally to the Embed function.
The worst-time complexities of FFG, with the dimensionality reduction, are summarized in Table IV.
FFG raw complexities adapted complexities

### v2

@@ IV-F1 Intensity and Impact of the Attack
To assess a plausible intensity of an attack that a graph can sustain without losing its utility, we show in Figure 4 a simple function performed on the three first graphs of Table II: the ranking of the top-100010001000 nodes according to the degree centrality [6]. The figure plots the changes in this ranking of the most central nodes, according to the classic Spearman metric [10]. It measures the correlation between two rankings of the top nodes (on the original graph and the attacked version). The correlation drops dramatically as the attack increases, even with a low value of 111 % of edges flipped. Thus, this random attack disrupts the connectivity of the central nodes, casting doubt on the usefulness of the resulting graph. In this context, 101010% of flipped edges is an extreme attack that appears to remove the utility of the graph, and can thus be considered a very conservative upper bound.
Figure 4: The effect of an edge flipping attack on three real graphs, regarding the degree centrality of the top-1000 most central nodes. The % of edges flips appears on the $x$-axis, while the Spearman correlation of the top-1000 ranking appears on the $y$-axis.
IV-F2 Robustness
Table II reports the performance of F&F on large real graphs. The key lengths align with those used by Eppstein et al. [13], while standard deviations produce watermarks with minimal distortion. Despite varying densities, all graphs received watermarks that caused only minor distortions, with edit distances ranging from $10^{-6}$ to $10^{-8}$, and reaching as low as $10^{-10}$ for the ”inf-belgium-osm” graph.
Graphs
@@ IV-F2 Robustness
$m$ $\sigma$ ED $\theta$ F&F
inf-belgium-osm
(1.4M, 1.1)
54 250 $3.10^{-10}$ 0.6
soc-YouTube-ASU
(1.1M, 2.6)
200 12k $1.1.10^{-8}$ 2
hollywood-2009
(1.1M, 52.7)
162 224k $1.6.10^{-8}$ 205
rgg-n-2-20-s0
(1M, 6.6)
119 1,624 $3.8.10^{-8}$ 10.2
kron-g500-logn20
(1M, 42.6)
71128 16k $3.5.10^{-8}$ 374
scale21-ef16-adj
(1.2M, 51)
44,743 12k $1.7.10^{-8}$ 233
roadNet-PA
(1.1M, 1.4)
48 184 $6.10^{-6}$ 0.6
delaunay-n20
(1M, 3.0)
63 61 $4.3.10^{-8}$ 6.2
TABLE II: Experiment settings and results for F&F, on real large graphs.
We now subject these graphs to attacks and observe the resilience of F&F. The last column shows that the extractions are successful despite the attack. We also note that additional tests have shown that all watermarks can also be extracted even under an attack that flips 100100100 % edges, for a value of $\theta$ set to be resistant to only 101010 % edge flips. It underscores the robustness of F&F, and the reach of Goal 4.
This experimental section confirms both the applicability of F&F to real graphs and the low distortion of the scheme on these graph structures, as well as its resilience to significant attack strengths.
@@ IV-F3 Densities and the Resulting Threshold
The final experiment for F&F is to test its robustness to different graph densities across the three graph models. We take the three graph generators (BA, WS, ER) and vary the density $\frac{|E|}{N}$ of graphs to 555, 303030, and 505050. 555 is a low density which respects the uniqueness condition, 303030 and 505050 are medium and high densities according to [26] and the graphs taken from SNAP [24]. For each generator and each density, we set $\theta$ to be resistant to $x$ edge flips ($x$-axis). This resilience is observed in 333 runs per point on Figure 5, where a dichotomous search sets $\theta$ to achieve the effective resilience to the attack.
Figure 5: The choice of $\theta$ in F&F to be resistant to $x$edge flips depending on the density and the type of graph model.
Figure 5 shows that the resulting threshold $\theta$ depends primarily on the type of graph. For BA or WS graphs, a setting $\theta=1$ is sufficient for F&F to be resistant to attacks with up to 101010% edge flips ($x$-axis), regardless of the density of the original graph. Density is the most relevant parameter when watermarking ER graphs, where $\theta$ can vary from 0.30.30.3 to 555 to be resistant to attacks of up to 10% edge flips.
@@ V A Comparative Benchmark
Finally, we perform a head-to-head comparative benchmark of F&F with the two state-of-the-art approaches [38] (Zhao) and [13] (Eppstein), both in terms of running time and resilience to attacks.
@@ V-A Related work
There are four steps in the graph embedding function of Zhao et al. [38]. First, they generate a random seed based on cryptographic keys. Then, they generate a random watermark graph (Erdős-Rényi) using the seed and match it with a subgraph of $G$. The selection of the subgraph relies on the structure of the vertices (i.e., their neighborhoods). Finally, an XOR operation performs the embedding between the subgraph and the random watermark graph. For the extraction, they regenerate the watermark graph and the subgraph, just as in the embedding process. After that, they identify in the potential $G^{*}$ which vertices are candidates to match the subgraph (again using the vertex structure). Finally, they extract the watermark using a recursive algorithm. Since the authors did not provide the code, we implemented it in Python 3. We note that our execution results are consistent with those presented by the authors in their experimental section (Figure 2b in [38]). We set $\theta=0.95$ and $B=25$, for an acceptable execution time.
@@
The scheme of Eppstein et al.[13] consists of the same three functions Keygen, Embed and Extract as the F&F scheme. In the scheme of Eppstein et al., private keys generated by the Keygen function are sets of vertex pairs in a graph containing as many vertices as in the graph to watermark. There are three main steps in the embedding function. The first step is to order and label the vertices of the graph, taking into account their degrees and using bit vectors. The vertices fall into three groups: high-degree, medium-degree, and low-degree vertices. Then, the procedure randomly flips each edge included in the private key according to the probability of its statistical existence in the graph. The function Extract uses graph isomorphism to identify the watermarked graph. Unlike F&F and Zhao [38], this method identifies a graph between multiple (simultaneously created) watermark copies, not just a single watermark graph with a given key. The selected copy is the closest copy considering the Hamming distance.
@@ V-B Execution Timings
We begin by benchmarking the runtime of the Embed and Extract functions from the competing watermarking schemes. Keys follow the generation guidelines from each paper, with timing results excluded because they execute very quickly. Implementations include scalability enhancements suggested by the original works.
@@
Figure 6: Benchmark comparison between Scheme, Eppstein [13], and Zhao [38]. (a) Timing ($y$-axis) of the Embed and Extract functions for the three competing schemes, depending on the BA graph size $N$ ($x$-axis). Executions taking less than the timeout (fixed at 25 minutes) are kept and plotted. (b) Robustness of the three schemes ($y$-axis) facing attacks: the normalized number of successful extractions, facing increasing edge flips (measured by edit distance on the $x$-axis as ratio of edges) on three large graphs.
In Figure 6(a), timings are measured in seconds as a function of the number of vertices $N$ in an increasing size BA graph (with an attachment parameter of 333). We set a timeout of 252525 minutes for the experiments (dotted gray horizontal line), so that all scheme executions– each running on a single core– are discarded if they exceed this threshold. For small and medium-sized graphs ($N\leq 10^{5}$), the embedding function of Zhao et al. is the most efficient method, while its extraction takes the longest. For medium and large graphs (more than $10^{5}$ vertices), the F&F scheme outperforms the state-of-the-art. Without exceeding the timeout, F&F can watermark graphs with 101010 times more vertices than the scheme of Eppstein et al. and up to 100100100 times more vertices than the scheme of Zhao et al.
Extractions take place on a non-attacked $G_{W}$, which constitutes the best-case scenario for the methods proposed by Eppstein et al. and Zhao et al.: their extractions are longer when applied to attacked graphs, because the approximate subgraph matching is more complex (while the matrix representation in F&F makes it time-invariant to attacks). The extractions of Zhao et al. and Eppstein et al. are linear in the number of vertices, while F&F is almost constant after $10k$ vertices.
@@ V-C Robustness to the Edge Flip Attack
Figure 6(b) shows the results for the three large graphs. The $y$-axis is the success rate. The $x$-axis shows the number of edges flipped, measured by the ED. Each experience averaged three times (the standard deviation is zero, as there is no variation in the results across all the performed executions).
We observe that F&F performs better than the scheme of Zhao et al. for the three graphs. The scheme of Zhao et al. fails before 101010% edges flipped, while F&F handles flips equivalent to 100100100% edges. F&F also beats the Eppstein scheme, except on the Flickr graph where Eppstein survives 100010001000% of flips.
We conclude that F&F is at least as robust– except on one configuration–, if not better, against random edge flips than its two competitors. Overall, the high robustness of F&F (resilience 100$\geq$ 100≥ 100% of flipped edges, see Table II and Figure 6b) makes it practical even in scenarios with extreme attack strengths.
@@ VI Conclusion
Large real graphs represent valuable assets that watermarking techniques can track for provenance. This paper bridges the gap between image processing and graph domains. While previous state-of-the-art schemes relied on standard graph manipulations, we have shown that knowledge and techniques from the multimedia community can be ported to the graph domain, yielding significant complexity improvements. The process involves studying the transformations between a graph and its corresponding real-valued image-like representation, and vice versa. Our method F&F shows strong performance on several metrics, including low false positives and negatives.
Future extensions of the framework may include other binarization methods, such as the optimized thresholding method proposed by Otsu [27]. It may lead to even smaller watermarks, which is crucial for minimizing distortions in the watermarked graph. Another promising direction is to adapt recent image watermarking schemes to our framework.
-A Notations
Table III is a summary of our main notations. In particular, the differences between $\omega$, $\mathcal{W}$ and $W$ are indicated.
@@
Figure 7: The F&F framework for watermarking unweighted graphs using an image domain scheme (hereby Cox et al. [4]).
-B Watermarking non-graph objects
@@
This paper is dedicated to watermark unweighted graphs. Recently, two graph watermarking approaches have been introduced that use purely graph techniques [13, 38] to watermark graphs. Since they are direct competitors to F&F, both schemes are described in detail in the main part of the paper (Section V), and then benchmarked.
@@
-C Theoretical Analysis of Uniqueness
@@
The only source of randomness in the scheme is the generation of the key; thus if $G$ is watermarked with another key $\omega_{2}$ resulting in $G_{W_{2}}$, then there is uniqueness if $G_{W_{1}}\neq G_{W_{2}}$. This property is not trivial because of the binarization which spreads keys over all the Fourier coefficients of the adjacency matrices. We then study the probability of uniqueness:
@@
Here the notation $A^{\prime}_{1,2}$ means that the result must be true for both $A^{\prime}_{1}$ and $A^{\prime}_{2}$ respectively and $av(A)$ is the average of all elements of the matrix $A$. By the linearity of the inverse Fourier transform, for $n\in\{1,2\}:A^{\prime}_{n}=A+FT^{-1}(\mathcal{W}_{n}).$
By construction, all the elements of each key $\omega$ (and its matrix version $\mathcal{W}$) are independent. The probability $P(G_{W_{1}}=G_{W_{2}})$ in Equation (1) can be computed with the cumulative distribution of $FT^{-1}(\mathcal{W}_{n})$ ($n\in\{1,2\}$).
First, note that: $FT^{-1}(\mathcal{W}_{n})=Z\mathcal{W}_{n}Z,$ where $Z[k,l]=\frac{1}{N}e^{2\imath\pi kl/N}$. Thus, $FT^{-1}(\mathcal{W}_{n})$ is a Gaussian matrix. With this, we can get an exact formula for the probability distribution function of this inverse Fourier transform flattened to one dimension [15]. In fact, let $\tilde{\mathcal{W}_{n}}$ be the flattened vector of $\mathcal{W}_{n}$, as $\forall k\in\{1,...,N^{2}\},\tilde{\mathcal{W}_{n}}[k]=\mathcal{W}_{n}[\lceil k/N\rceil,(k-1)\mod N+1],$ where $\lceil k/N\rceil$ being the upper integer part of $k/N$ and $(k-1)\mod N$ being the rest of the Euclidean division of $k-1$ by $N$. Using these notations, we easily show that: $FT^{-1}(\tilde{\mathcal{W}_{n}})=\tilde{Z}\mathcal{W}_{n},$ where $\tilde{Z}=Z\bigotimes Z$ is the $N^{2}\times N^{2}$ Kronecker product of $Z$ with itself. Let $(z_{k,l})$ be the coefficients of $Z$, then for $(k^{\prime},l^{\prime})\in\{1,...,N^{2}\}$: $(\tilde{Z})_{k^{\prime},l^{\prime}}=z_{\lceil k^{\prime}/N\rceil,\lceil l^{\prime}/N\rceil}z_{(k^{\prime}-1)\mod N+1,(l^{\prime}-1)\mod N+1}.$ Since $\tilde{\mathcal{W}_{n}}$ is a Gaussian vector, $FT^{-1}(\tilde{\mathcal{W}_{n}})$ is also a Gaussian vector. Its average and standard deviation are computed using the matrix $\tilde{Z}$. In conclusion, the watermark’s uniqueness is calculable.
@@
The connection between graph densities and collisions is given by the theoretical analysis in Equation (-C). Indeed, as $av(A)=\frac{|E|}{N^{2}}$, the threshold for binarization is equal to the density divided by $N$. For the $N=1M$ graphs (density less than 101010) in Figure 2 and the given keys: $av(A)\simeq 0,$ $av(|FT^{-1}(W)|)\simeq 0,$ and $av(|A+FT^{-1}(W)[i,j]|)\simeq av(A).$
Then, it is clear that the denser a graph, the more $A[i,j]$ coefficients are equal to 111, and decreases the probability of collision (as seen in Equation (-C)).
Successful Embeddings: a Corollary of Uniqueness
In this section, we show that studying a successful embedding is a corollary of the analytical form of the uniqueness. The embedding of a key $\omega$ in a graph $G$ is successful if the watermarked graph $G_{W}$ is different from the original graph $G$. It can happen that the binarization discards a key built with insufficiently large parameters, so that no watermark is inserted into the graph. The study of successful embedding is therefore the estimation of $P(G_{W}=G)$.
@@
[…]
Note that in practice, the owner of the graph knows when the embedding has succeeded or failed: if the watermarked graph is equal to the original graph after an embedding attempt (which is linear to check in $O(|E|)$), it is sufficient for the owner to retry an embedding operation with larger key parameters until the two graphs are truly different.
-D Scalability reduction
In Section IV-B1, we explain that a scability reduction is applied to F&F as it is the case in  [13, 38]. This section provides the details and explanation of such an operation.
When considering graphs, the most significant information is contained in the vertices of the highest degree vertices; thus, analogous to low frequencies, in F&F, the watermark is inserted into the lowest amplitude coefficients in the spectrum among the higher-degree vertices. While the size of images on which the Cox et al. scheme is typically applied is in the order of tens of thousands of pixels, we also choose to take the top $N_{0}=10^{4}$ higher-degree vertices of $G$ to apply F&F.
In this way, Embed and Extract are applied only on $N_{0}$ vertices and not on the $N$ vertices of the original graph. Their worst-case time complexities, computed in Section III-D, are now $\mathcal{O}\left(N_{0}^{2}\log N_{0}\right)$. However, to determine the subgraph on which to apply F&F, it is necessary to select the $N_{0}$ vertices of highest degrees (which is calculated in $\mathcal{O}\left(N\log N_{0}\right)$). This cost is added naturally to the Embed function.
The worst-time complexities of F&F, with the dimensionality reduction, are summarized in Table IV.
F&F raw complexities adapted complexities